### v1

@@ Abstract
Software bugs cost the global economy billions of dollars annually and claim $\sim$50of the programming time from software developers. Locating these bugs is crucial for their resolution but challenging. It is even more challenging in deep-learning systems due to their black-box nature. Bugs in these systems are also hidden not only in the code but also in the models and training data, which might make traditional debugging methods less effective. In this article, we conduct a large-scale empirical study to better understand the challenges of localizing bugs in deep-learning systems. First, we determine the bug localization performance of four existing techniques using 2,365 bugs from deep-learning systems and 2,913 from traditional software. We found these techniques significantly underperform in localizing deep-learning system bugs. Second, we evaluate how different bug types in deep learning systems impact bug localization. We found that the effectiveness of localization techniques varies with bug type due to their unique challenges. For example, tensor bugs were more accessible to locate due to their structural nature, while all techniques struggled with GPU bugs due to their external dependencies. Third, we investigate the impact of bugs’ extrinsic nature on localization in deep-learning systems. We found that deep learning bugs are often extrinsic and thus connected to artifacts other than source code (e.g., GPU, training data), contributing to the poor performance of existing localization methods.
@@ 1 Introduction
To date, there exist only a few techniques for detecting bugs in deep learning systems. Most of them concentrate on specific type of bugs (e.g., model bugs, training bugs) without considering the broader spectrum of deep learning systems. Wardat et al. [18] propose a dynamic approach to localize different types of model bugs in the Deep Neural Network (DNN). They identify the faulty layers containing numerical bugs by customizing the Keras’ callback function and analyzing the dynamic behaviors of a model. However, their solution focuses on only model bugs from deep learning systems, strongly coupled with the Keras library, and achieves a low accuracy, which presents significant challenges for a widespread adoption by the industry. In another study, Wardat et al. [20] propose a heuristic-based approach to diagnose two main categories of bugs – model bugs and training bugs. They also recommend actionable fixes of the bugs based on the diagnosis. Since their approach depends on a set of hard-coded rules, it might be limited in terms of scalability and context-awareness. In a recent work, Cao et al. [21] introduce a technique that leverages the dynamic properties of a model and an ensemble of three machine learning classifiers (e.g., KNN, Decision Tree, Random Forest) to localize five types of training bugs (e.g., loss, gradient) from deep learning systems. Their technique might also not be able to address a broader array of bug types from deep learning systems, highlighting the issues of scalability. On the other hand, Kim et al. [22] use basic Information Retrieval (IR) algorithms, such as rVSM and BM25, to localize bugs in deep-learning systems. They report poor performance but do not perform any comprehensive analysis to understand the poor performance of IR-based techniques.
@@
In this article, we conduct a large-scale empirical study to better understand the challenges of locating bugs in deep learning systems. First, we collect a total of 2,365 bugs from deep-learning systems and 2,913 bugs from traditional software systems (a.k.a non-deep-learning systems), and empirically show how existing techniques (e.g., BugLocator [11], BLUiR [12], BLIA [23], DNNLOC [17]) perform in locating bugs from deep learning systems. Our work utilizes these traditional techniques as a foundational framework, adapting their core principles to the specific nuances of deep learning bugs. Second, we categorize our collected bugs based on an existing bug taxonomy [30] and found that certain bugs from deep learning systems (e.g., GPU bugs) are more difficult than others to locate due to their multifaceted heterogeneous dependency issues. Finally, we found that deep learning bugs are connected to artifacts other than source code (e.g., GPU, training data, external dependencies) and are prone to be extrinsic in nature, which might explain the poor performance of existing techniques for these bugs. We thus answer three important research questions in our study as follows.
@@
[…]
    We evaluated the performance of four existing approaches (BugLocator [11], BLUiR [12], BLIA [23], and DNNLOC [17]) using two datasets – Denchmark [31] and BugGL [32]. First, we found that their performance measures are poorer (e.g., 31.59% less MAP for BugLocator, 33.25% for BLUiR, 34.14% for BLIA, 31.43% for DNNLOC) in localizing bugs from deep learning systems than that of non-deep learning systems. Our statistical tests (t-test [33], Cohen’s D [34]) also report that their performance is significantly lower. Second, we found that localizing bugs from the deep learning frameworks is more challenging than libraries or tools due to the frameworks’ inherent complexity. Although our findings reinforce the existing understanding and belief about the challenges of the bugs in deep learning systems [35, 22, 18], we also substantiate them with solid empirical evidence and demonstrate the performance gap of existing solutions in localizing the two categories of bugs.
@@
[…]
    We use an existing taxonomy [30] of bugs to classify the bugs in deep learning systems and evaluate the performance of four existing techniques for each type of bug. First, we found that 64.80% of the bugs from deep learning systems are related to deep learning (e.g., model, training), whereas the remaining ones are not. Second, we found that DNNLOC demonstrated better results in locating model and tensor bugs, possibly due to its ability to capture comprehensive contextual information specific to these bugs. We also found that BLUiR performs comparably to DNNLOC for training bugs, which might be attributed to its structured information retrieval. However, all four baseline techniques experienced difficulty localizing GPU bugs. Thus, our analysis offers valuable insights regarding the nature of different types of bugs in deep learning systems and highlights the specific strengths and weaknesses of existing techniques, which could be useful to advance debugging support for deep learning systems.
@@
[…]
    Bugs triggered by external entities (e.g., third-party libraries, GPU) are called extrinsic bugs [36]. Given the frequent use of deep learning libraries and their external dependencies, the bugs in deep learning systems could be extrinsic [10]. Since the existing techniques mostly focus on intrinsic bugs (i.e., triggered by bug-introducing change), we investigate how they deal with extrinsic bugs from deep learning systems. First, we found deep learning systems have 40.00% extrinsic bugs, which is almost four times higher than that of non-deep learning systems. Second, we found that the localization performance of existing techniques degrades significantly for extrinsic bugs (e.g., 15.20% less MAP for DNNLOC) (Table 21). We also found that deep neural network-based solution (e.g., DNNLOC) is not particularly helpful for locating extrinsic bugs either because they are designed to detect code patterns, not issues from external sources or environments. Finally, we found a significant correlation between the bugs in deep learning systems and the extrinsic factors using appropriate statistical analysis (Chi-Square test), which delivers valuable insight for designing effective solutions to find bugs in deep learning systems.
@@ 2 Background
In this section, we introduce the necessary terminologies and concepts to follow the remainder of the article. We introduce extrinsic bugs, intrinsic bugs, and the taxonomy of deep-learning bugs.
2.1 Extrinsic bug
A bug caused by the factors external to a software system, such as changes to the operating environment, requirements, or third-party libraries, is known as extrinsic bug. Rodriguez-Perez et al. suggest three heuristics based on bug reports to identify extrinsic bugs as follows [36].
(b) Requirement: An extrinsic bug is triggered by a change outside of the project’s version control system. During software development, if a user requirement gets changed after implementation, the development team might implement the new requirement without discarding the old feature. The old, unexpected feature will then be considered as an extrinsic bug.
2.2 Intrinsic bug
@@ 2.3 Taxonomy of bugs in deep learning systems
  • GPU bug is connected to the Graphics Processing Unit (GPU) used in the system (Table 18 [43]). For example, if the model’s memory requirements exceed the available GPU memory, or the GPU is not compatible with the DL framework, then they could lead to errors during model training.
@@
As shown in Table 1, Bug 1860 [37] is a deep learning-related extrinsic bug triggered by the change in the environment. When the WMT19 model runs on multiple GPUs, the execution fails since the same GPU cannot store both the model and data. It is clearly related to the deep learning module. On the other hand, this bug is not related to the Fairseq library (a.k.a., deep learning application), rather, it is related to external factors (e.g., GPU), which indicates its extrinsic nature.
@@ 3 Study Methodology
Fig. 1 shows the schematic diagram of our conducted study. First, we collect bug reports from two benchmark datasets for two different software systems: deep learning systems [31] and traditional software systems [32]. Then, we contrast the performance of four existing techniques [11, 12, 23, 17] in locating bugs between deep learning systems [31] and traditional software systems [32]. Second, we perform an in-depth analysis to understand the challenges of localizing different types of deep-learning bugs. Finally, we investigate the influence of extrinsic factors on deep learning bugs and their impact on bug localization. This section discusses the major steps of our study design as follows.
@@ 3.1 Construction of dataset
Dataset collection. In our study, we use two benchmark datasets – BugGL and Denchmark – that have been previously used by the literature [32, 22]. BugGL [32] contains bugs from Python-based, traditional software systems (a.k.a non-deep learning-based systems), whereas Denchmark focuses on bugs from deep learning-based systems. BugGL contains a total of 2,913 bug reports from 12 Python projects [32]. On the other hand, the original Denchmark dataset [31] contains 4,577 bug reports from 193 deep learning-based projects, which are written in ten programming languages (JavaScript, Python, Java, Go, C++, Ruby, TypeScript, PHP, C#, and C). We used 2,365 bug reports from 136 deep learning-based projects (written in Python) from the Denchmark dataset [22]. We limited our dataset to Python-based bugs to ensure a fair contrastive analysis between deep learning bugs and traditional software bugs.
@@
We adopted a set of filtration steps to construct our experimental dataset, which consists of 86 projects from DL systems (shown in Table 2). First, the initial Denchmark dataset has 136 projects across various classes, including client-based applications, frameworks, tools, libraries, search engines, and compilers. Our study focuses on core elements of deep learning – frameworks (e.g., Apache MXNet), libraries (e.g., OpenCV), and tools (e.g., Fairseq) – rather than application-specific software (e.g., PhotoPrism). We thus collect the systems pertaining to those three categories and discard all application-specific systems. Second, we checked for overlapping projects among these classes to ensure distinct categorization. Finally, from a total of 136 projects, we finalized our experimental dataset with 17 projects categorized under frameworks, 28 under tools, and 41 under libraries, bringing the total to 86 projects from deep learning systems.
Deep learning systems differ significantly from traditional software systems due to the complexity of model integration, intricate interactions between deep learning libraries [10], and multifaceted dependencies.
@@
Corpus creation. We first download the latest version of the code repositories, ensuring that we have the most up-to-date code for analysis. Next, to accurately link bug reports to their corresponding buggy code, we used the heuristic of Kim et al. [22], focusing on commit messages and bug reports. This involved analyzing commit messages for keywords indicative of bug fixes (e.g., ‘fix’, ‘bug’, ‘error’, ‘resolve’) and connecting them to corresponding bug reports. To identify the correct bug-fix commit, we cross-referenced these commit messages with bug IDs from the reports, ensuring a precise match. To extract the buggy part from the bug-fix commit, we employed PyDriller, suggested by Kim et al. [22]. We also capture each project’s most recently released version as of the bug report’s date, and collect the appropriate version of the buggy code, especially in cases where the bug reports did not have clear buggy version information.
Query construction. In IR-based bug localization, bug reports are treated as queries that can be executed with a search engine to detect the relevant source documents from the corpus. We construct a repository of bug reports by parsing the original datasets (Denchmark & BugGL) and extracting important information such as bug IDs, descriptions, and timestamps. We construct the queries by extracting tokens from the title and description of bug reports, removing stop words, stemming each word, and splitting the tokens.
@@
Categorizing Classes in Deep Learning Systems. In the original dataset by Kim et al. [31], they have identified different categories (e.g., frameworks, libraries, tools) in deep learning systems. However, there were some projects that overlapped in terms of being frameworks, libraries, or tools. This is because the lines between these categories might be unclear sometimes. Some libraries may evolve into frameworks or vice versa, and tools could be integrated into either libraries or frameworks. To address this issue, we removed the overlapping projects from our dataset. We then manually analyzed each selected project based on the official documentation to ensure they clearly belonged to one of the three categories. This process helped us maintain distinct differences among our study’s frameworks, libraries, and tools. We have spent $\approx$5 hours on the manual analysis. We have provided the list of the selected projects from these three categories, along with their detailed project descriptions, in our replication package [46].
@@
Data #P #BR #SF #BF #Versions Mean Max Mean Max Mean Max Dench. 136 1795 441.60 3,559 2.60 227 6.7 80 BugGL 12 1795 420.50 3,306 2.35 198 6.2 65 P: Projects, BR: Bug Reports, SF: Source Files, BF: Buggy Files

### v2

@@ Abstract
Software bugs cost the global economy billions of dollars annually and claim $\sim$50of the programming time from software developers. Locating these bugs is crucial for their resolution but challenging. It is even more challenging in deep-learning systems due to their black-box nature. Bugs in these systems are also hidden not only in the code but also in the models and training data, which might make traditional debugging methods less effective. In this article, we conduct a large-scale empirical study to better understand the challenges of localizing bugs in deep-learning systems. First, we determine the bug localization performance of five existing techniques using 2,365 bugs from deep-learning systems and 2,913 from traditional software. We found these techniques significantly underperform in localizing deep-learning system bugs. Second, we evaluate how different bug types in deep learning systems impact bug localization. We found that the effectiveness of localization techniques varies with bug type due to their unique challenges. For example, tensor bugs were more accessible to locate due to their structural nature, while all techniques struggled with GPU bugs due to their external dependencies. Third, we investigate the impact of bugs’ extrinsic nature on localization in deep-learning systems. We found that deep learning bugs are often extrinsic and thus connected to artifacts other than source code (e.g., GPU, training data), contributing to the poor performance of existing localization methods.
@@ 1 Introduction
To date, there exist only a few techniques for detecting bugs in deep learning systems. Most of them concentrate on specific types of bugs (e.g., model bugs, training bugs) without considering the broader spectrum of deep learning systems. Wardat et al. [18] propose a dynamic approach to localize different types of model bugs in the Deep Neural Network (DNN). They identify the faulty layers containing numerical bugs by customizing the Keras’ callback function and analyzing the dynamic behaviors of a model. However, their solution focuses on only model bugs from deep learning systems, strongly coupled with the Keras library, and achieves a low accuracy, which presents significant challenges for widespread adoption by the industry. In another study, Wardat et al. [20] propose a heuristic-based approach to diagnose two main categories of bugs – model bugs and training bugs. They also recommend actionable fixes of the bugs based on the diagnosis. Since their approach depends on a set of hard-coded rules, it might be limited in terms of scalability and context-awareness. In a recent work, Cao et al. [21] introduces a technique that leverages the dynamic properties of a model and an ensemble of three machine learning classifiers (e.g., KNN, Decision Tree, Random Forest) to localize five types of training bugs (e.g., loss, gradient) from deep learning systems. Their technique might also not be able to address a broader array of bug types from deep learning systems, highlighting the issues of scalability. On the other hand, Kim et al. [22] use basic Information Retrieval (IR) algorithms, such as rVSM and BM25, to localize bugs in deep-learning systems. They report poor performance but do not perform any comprehensive analysis to understand the poor performance of IR-based techniques.
@@
In this article, we conduct a large-scale empirical study to better understand the challenges of locating bugs in deep learning systems. First, we collect a total of 2,365 bugs from deep-learning systems and 2,913 bugs from traditional software systems (a.k.a non-deep-learning systems), and empirically show how existing techniques (e.g., BugLocator [11], BLUiR [12], BLIA [23], DNNLOC [17]), bjXnet [30]) perform in locating bugs from deep learning systems. Our work utilizes these traditional techniques as baselines, adapting their core principles to the specific nuances of deep learning bugs. Second, we categorize our collected bugs based on an existing bug taxonomy [31] and found that certain bugs from deep learning systems (e.g., GPU bugs) are more difficult than others to locate due to their multifaceted heterogeneous dependency issues. Finally, we found that deep learning bugs are connected to artifacts other than source code (e.g., GPU, training data, external dependencies) and are prone to be extrinsic in nature, which might explain the poor performance of existing techniques for these bugs. We thus answer three important research questions in our study as follows.
@@
    RQ1: How effective are the existing approaches in localizing bugs from deep learning systems?
    We evaluated the performance of five existing approaches (BugLocator [11], BLUiR [12], BLIA [23], DNNLOC [17], and bjXnet [30]) using two datasets – Denchmark [32] and BuGL [33]. First, we found that their performance measures are poorer (e.g., 31.59% less MAP for BugLocator, 33.25% for BLUiR, 34.14% for BLIA, 31.43% for DNNLOC, 27.87% for bjXnet ) in localizing bugs from deep learning systems than that of non-deep learning systems. Our statistical tests (t-test [34], Cohen’s D [35]) also report that their performance is significantly lower. Second, we found that localizing bugs from the deep learning frameworks is more challenging than libraries or tools due to the frameworks’ inherent complexity. Although our findings reinforce the existing understanding and belief about the challenges of the bugs in deep learning systems [36, 22, 18], we also substantiate them with solid empirical evidence and demonstrate the performance gap of existing solutions in localizing the two categories of bugs.
@@
    RQ2: How do different types of bugs in deep learning systems impact bug localization?
    We use an existing taxonomy [31] of bugs to classify the bugs in deep learning systems and evaluate the performance of five existing techniques for each type of bug. First, we found that 64.80% of the bugs from deep learning systems are related to deep learning (e.g., model, training), whereas the remaining ones are not. Second, we found that bjXnet and DNNLOC demonstrated better results in locating model and tensor bugs, possibly due to their ability to capture comprehensive contextual information specific to these bugs using deep neural networks. We also found that BLUiR performs comparably to DNNLOC for training bugs, which might be attributed to its structured information retrieval. However, all five baseline techniques experienced significant difficulty localizing GPU bugs. Thus, our analysis offers valuable insights regarding the nature of different types of bugs in deep learning systems and highlights the specific strengths and weaknesses of existing techniques, which could be useful to advance debugging support for deep learning systems.
@@
    RQ3: What are the implications of extrinsic bugs in deep learning systems for bug localization?
    Bugs triggered by external entities (e.g., third-party libraries, GPU) are called extrinsic bugs [37]. Given the frequent use of deep learning libraries and their external dependencies, the bugs in deep learning systems could be extrinsic [10]. Since the existing techniques mostly focus on intrinsic bugs (i.e., triggered by bug-introducing change), we investigate how they deal with extrinsic bugs from deep learning systems. First, we found deep learning systems have 40.00% extrinsic bugs, which is almost four times higher than that of non-deep learning systems. Second, we found that the localization performance of existing techniques degrades significantly for extrinsic bugs (e.g., 15.20% less MAP for DNNLOC) (Table 30). We also found that deep neural network-based solutions (e.g., DNNLOC) are not particularly helpful for locating extrinsic bugs because they are designed to detect code patterns, not issues from external sources or environments. Finally, we found a significant correlation between the bugs in deep learning systems and the extrinsic factors using appropriate statistical analysis (Chi-Square test), which delivers valuable insight for designing effective solutions to find bugs in deep learning systems.
@@ 2 Background
In this section, we introduce the necessary terminologies and concepts to follow the remainder of the article. We introduce extrinsic bugs, intrinsic bugs, the taxonomy of deep-learning bugs, and the categories of projects in deep-learning systems.
@@
2.1 Extrinsic bug
A bug caused by factors external to a software system, such as changes to the operating environment, requirements, or third-party libraries, is known as extrinsic bug. Rodriguez-Perez et al. suggests three heuristics based on bug reports to identify extrinsic bugs as follows [37].
(b) Requirement: An extrinsic bug is triggered by a change outside the project’s version control system. During software development, if a user requirement changes after implementation, the development team might implement the new requirement without discarding the old feature. The old, unexpected feature will then be considered an extrinsic bug.
2.2 Intrinsic bug
@@ 2.3 Taxonomy of bugs in deep learning systems
  • GPU bug is connected to the Graphics Processing Unit (GPU) used in the system (Table 27 [44]). For example, if the model’s memory requirements exceed the available GPU memory or the GPU is not compatible with the DL framework, then they could lead to errors during model training.
@@
As shown in Table 1, Bug 1860 [38] is a deep learning-related extrinsic bug triggered by the change in the environment. When the WMT19 model runs on multiple GPUs, the execution fails since the same GPU cannot store both the model and data. It is clearly related to the deep learning module. On the other hand, this bug is not related to the Fairseq library (a.k.a., deep learning application); rather, it is related to external factors (e.g., GPU), which indicates its extrinsic nature.
@@
2.4 Categories of projects in deep learning systems
Deep learning-based projects can be categorized into Frameworks, Tools, Libraries, Applications, Engines, Platforms, and Compilers [32]. The core elements—Frameworks, Tools, and Libraries—comprise the majority of relevant bug reports, accounting for 96.6% of the dataset [32].
Table 2: Example of a bug report from deep learning framework [45]
Framework Bug (Bug ID: 10224)
Title
Language model example cannot be run
Description
The language model example cannot be run without manually
creating a data folder. There are also inconsistencies between the
documentation and the code. Optional argument – data DATA
location of the data corpus does not appear in the code train.py
Detailed BR: https://github.com/apache/mxnet/issues/10224
A deep learning framework is a software platform that provides the environment for designing, training, and deploying deep learning models [46]. Examples include TensorFlow [^1] https://www.tensorflow.org/, PyTorch [^2] https://pytorch.org/, and Apache MXNet [^3] https://mxnet.apache.org/versions/1.9.1/. These frameworks come with pre-defined modules and functions and offer a structured way to implement deep learning architectures using high-level programming interfaces [47]. The example bug in Table 2 [45] is characterized by the inability to run a language model without manually creating a data folder. It represents a framework bug because it directly impacts the core functionalities of the framework, specifically the design and training of models. Frameworks are expected to provide seamless, user-friendly environments for developing deep learning models, and issues that hinder the ease of use, such as documentation inaccuracies and additional manual setup steps, indicate problems at the framework level.
Table 3: Example of a bug from deep learning library [40]
Library Bug (Bug ID: 313)
Title
A bug in GPT2Tokenizer
Description
GPT2Tokenizer fails to recover a sentence
\”BART is a seq2seq model.\”
with encoded ids of it.
The output sentence is \”BART is a seqseq model.\”.
It should be related to numbers’ processing.
URL: https://github.com/tanyuqian/texar-pytorch
/blob/master/examples/bart/gpt2_tokenizer_bug.py
Table 4: Example of a bug from deep learning tool [49]
Tool Bug (Bug ID: 5596)
Title
Tensorboard can not load all Hyperparameters keys
Description
Encountered an issue in TensorBoard where it couldn’t load
all hyperparameter keys when I used writer.add-hparams with
different hparam-dict parameters in multiple experiments. This
problem made it difficult to track and compare different
hyperparameter settings across these experiments, affecting the
overall functionality of TensorBoard. I provided code snippets and
responses to showcase the inconsistency in the display of
hyperparameters in TensorBoard’s interface. This issue
is significant as it hinders the effective use of TensorBoard
for experiment tracking and analysis.
@@ 3 Study Methodology
Fig. 1 shows the schematic diagram of our conducted study. First, we collect bug reports from two benchmark datasets for two different software systems: deep learning systems [32] and traditional software systems [33]. Then, we contrast the performance of five existing techniques [11, 12, 23, 17] in locating bugs between deep learning systems [32] and traditional software systems [33]. Second, we perform an in-depth analysis to understand the challenges of localizing different deep-learning bugs. Finally, we investigate the influence of extrinsic factors on deep learning bugs and their impact on bug localization. This section discusses the major steps of our study design, which are as follows:
@@ 3.1 Construction of dataset
Dataset Collection. In our study, we use two benchmark datasets: BuGL and Denchmark. BuGL contains traditional software bugs, while Denchmark focuses on bugs in deep learning systems. BuGL [33] provides a wide range of traditional software bugs across multiple programming languages (e.g., C++, Java, Python) and includes a total of 2,913 bug reports from 12 Python projects. In contrast, Denchmark [32], published in the MSR 2021 Dataset Track, focuses on bugs in deep learning systems written in ten different programming languages, including JavaScript, Python, Java, Go, C++, Ruby, TypeScript, PHP, C#, and C. It contains 2,365 bug reports from 136 deep learning-based projects written in Python. While the Denchmark dataset may not cover all possible systems, it is representative of the current state of deep learning system debugging, encompassing a wide range of 136 DL projects (e.g., computer vision, natural language processing, image processing, and visualization).
To ensure a fair comparison between deep learning and traditional software bugs, we only included Python-based projects and bugs from both datasets. Additionally, multiple studies [33, 22] have used Denchmark and BuGL for bug localization experiments, making them suitable benchmarks for our experiments as well.
@@
We adopted a set of filtration steps to construct our experimental dataset, which consists of 86 projects from DL systems (shown in Table 5).
  • Original Denchmark Dataset: The original Denchmark dataset contains 136 projects and 2,365 bug reports. These projects are divided into several categories by the original authors [32]: Frameworks, Tools, Libraries, Applications, Engines, Platforms, and Compilers.
  • Filter 1: We focused on the core elements of deep learning – Frameworks, Tools, and Libraries, as they represent the majority of relevant bug reports in the dataset (96.6%). In the original dataset, there were 25 projects from Framework, 31 projects from Tool, and 44 projects, which gave us a total of 100 projects. We excluded projects with fewer than ten reported bugs to ensure a robust dataset for meaningful comparisons of bug localization techniques. After applying this filter, we had 18 framework projects, 29 tool projects, and 42 library projects, which led to a total of 87 projects.
  • Filter 2: We noted that some projects had multiple categories (e.g., a tool that also functions as a library). After carefully reviewing the official documentation and analyzing the role of these projects, we removed overlapping projects to ensure that each project was uniquely classified as either a framework, tool, or library.
After applying these filtering steps, the final experimental dataset consists of 86 projects and 1,795 bug reports - Frameworks: 17 projects, 746 bugs, Tools: 28 projects, 455 bugs, Libraries: 41 projects, 594 bugs.
@@
Corpus creation. We first download the latest version of the code repositories, ensuring that we have the most up-to-date code for analysis. Next, to accurately link bug reports to their corresponding buggy code, we used the heuristic of Kim et al. [22], focusing on commit messages and bug reports. This involved analyzing commit messages for keywords indicative of bug fixes (e.g., ‘fix’, ‘bug’, ‘error’, ‘resolve’) and connecting them to corresponding bug reports. To identify the correct bug-fix commit, we cross-referenced these commit messages with bug IDs from the reports, ensuring a precise match. To extract the buggy part from the bug-fix commit, we employed PyDriller, suggested by Kim et al. [22]. We also capture each project’s most recently released version as of the bug report’s date and collect the appropriate version of the buggy code, especially in cases where the bug reports did not have clear buggy version information.
Query construction. In IR-based bug localization, bug reports are treated as queries that can be executed with a search engine to detect the relevant source documents from the corpus. We construct a repository of bug reports by parsing the original datasets (Denchmark & BuGL) and extracting important information such as bug IDs, descriptions, and timestamps. We construct the queries by extracting tokens from the title and description of bug reports, removing stop words, stemming each word, and splitting the tokens.
@@
Data #P #BR #SF #BF #Versions Mean Max Mean Max Mean Max Dench. 136 1795 441.60 3,559 2.60 227 6.7 80 BuGL 12 1795 420.50 3,306 2.35 198 6.2 65 P: Projects, BR: Bug Reports, SF: Source Files, BF: Buggy Files